OpenStudy (anonymous):

write this in a + bi form: 4/3i +2/5i

OpenStudy (jdoe0001):

what would be the LCD?

OpenStudy (anonymous):

15i right?

OpenStudy (jdoe0001):

yeap thus one sec

OpenStudy (anonymous):

sure cx

OpenStudy (jdoe0001):

\(\bf \cfrac{4}{3i}+\cfrac{2}{5i}\implies \cfrac{(4\cdot 5)+(3\cdot 2)}{15i}\implies \cfrac{(20)+(6)}{15i}\implies \cfrac{26}{15i} \\ \quad \\ \textit{now, to rationalize it, we multiply top and bottom by }i \\ \quad \\ \cfrac{26}{15i}\cdot \cfrac{i}{i}\implies \cfrac{26i}{15i^2}\qquad {\color{brown}{ i^2=-1}}\qquad \cfrac{26i}{15({\color{brown}{ -1}})}\implies \cfrac{26i}{-15}\implies ?\)

OpenStudy (anonymous):

dont you multiply -i?

OpenStudy (jdoe0001):

nope.... -i? to "rationalizate", we would want to get rid of the "i" in the denominator so we use "i" to get \(i^2\)

OpenStudy (jdoe0001):

\(\bf \cfrac{26}{15i}\cdot \cfrac{i}{i}\implies \cfrac{26i}{15i^2}\qquad {\color{brown}{ i^2=-1}}\qquad \cfrac{26i}{15({\color{brown}{ -1}})}\implies \cfrac{26i}{-15} \\ \quad \\ -\cfrac{26i}{15}\implies \begin{cases} a&+bi\\ \bf 0&-\cfrac{26}{15}i \end{cases}\)

OpenStudy (anonymous):

oh ok thnx cx